OpenStudy (snowcrystal):

Help please I dont understand this...I will fan and medal!!!!! A rope is swinging in such a way that the length of the arc is decreasing geometrically. If the 1st arc is 18ft long and the 3rd arc is 8ft long, what is the length of the 2nd arc? A.) 12ft B.) 10ft C.) 5ft D.) 72ft

OpenStudy (snowcrystal):

@K12awesomeness

OpenStudy (kitty626oldrose):

b

OpenStudy (k12awesomeness):

Hold on...

OpenStudy (k12awesomeness):

It's A.. 12, because the second rope be x..then x/18 = 8/x or x^2 = 18 x 8 = 144 x =12...this is the answer

OpenStudy (k12awesomeness):

Hope this helped:)

OpenStudy (snowcrystal):

thank you :3 ^_^

OpenStudy (k12awesomeness):

No Problem!! tag me if u need me

OpenStudy (snowcrystal):

i will